Fermi National Laboratory


Accelerator Update for June 3, 2005 - June 6, 2005

Friday June 3
The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and Meson MCenter experiments taking beam.

Operations made the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls ready for access at 8:19 AM. CDF finished their access at 9:40 AM.

Operations established beam to the Meson MTest experiment at 10:48 AM.

NuMI experts began conducting target scans at 11:14 AM.

D0 came out of access and operators began turning the Tevatron (TeV) on at 1:06 PM.

Operations put the TeV into a dry squeeze at 4:03 PM.

Operations made D0 ready for an access at 4:36 PM.

D0 came out of access at 11:35 PM.

MTest experimenters reported at 11:52 PM that they were finished with beam for the night. An operator disabled the beam.

Saturday June 4
The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the TeV through a wet squeeze at 1:31 AM.

Operations put the TeV into shot setup around 2:05 AM.

At 2:28 AM, an Accumulator abort kicker (AAK1) wouldn't fire properly, nor would it pull the abort. This problem put shot setup on hold. Operators called in two Pbar kicker experts.

The Pbar kicker experts reported at 4:25 AM that they found and repaired a bad fiber optic cable.

Operations restarted the TeV shot setup process at 5:04 AM.

At 5:46 AM, an operator reported that she was beginning to load the TeV with protons, but at 6:03 AM, a problem with Booster intensity was delaying the process. Operations called in a Booster expert.

Operations reported around 8 AM that they had started loading the final protons.

At 9:01 AM, the TeV suffered a sector F3 quench. Investigations discovered a sequencer problem that allowed beam to be injected on top of other beam.

Operations reestablished beam to NuMI at 9:39 AM.

Operations put the TeV through a dry squeeze at 11:49 AM.

Operations put the TeV into shot setup around 12:50 PM.

At 1:32 PM, Operations reported an approaching thunderstorm. It passed by with few problems to the shot setup. The village did lose power for a short time.

Operations established combination store 4193 at 5:33 PM with an initial luminosity of 90.58E30.

NuMI resumed taking beam at 6:04 PM. Pbar resumed stacking at 6:06 PM.

Operations resumed sending beam to MCenter at 6:22 PM. A SWIC failure wouldn't allow operators to view the beam. Operators rebooted the SWICS, which brought back some of them. The rest of the SWICS were operational by the end of the shift.

At 10:45 PM, Operations began preparing Booster for an access to replace a Klixon. Booster resumed accelerating beam by 11:52 PM.

Sunday June 5
The midnight shift began with Operations monitoring store 4193, with Pbar stacking, and with beam to MiniBooNE, NuMI, and MCenter.

Operations put the Recycler into shot setup to stash antiprotons at 1:13 AM. The Recycler finished stashing by 2:52 AM.

At 5:43 AM, Operations reported a Booster RF station (BRF2) tripped off and wouldn't reset. Operations contacted experts.

Operations prepared the Booster for an access to investigate the BRF2 problem. The experts had the problem fixed an hour later.

Pbar resumed stacking at 8:16 AM and resumed sending beam to MiniBooNE, NuMI, and MCenter a few minutes later.

At 10:39 AM, Operation reported that they had been monitoring a Pbar LCW leak. Operators had checked all the service buildings and not found it, which meant the leak was in the tunnel. Operations began preparing the Recycler for a stash at 10:46 AM so that the Pbar tunnels could be accessed.

The Recycler stash ended around noon. Operations began preparing Pbar for an access.

At 1:05 PM, the Pbar access team reported that the LCW leak appeared to be in the transport line. This access shut off beam to MiniBooNE, NuMI, and MCenter. They found a leaking hose on a quadrupole magnet (IQ18).

At 1:36 PM, Operations reported that the Pbar interlocks were accidentally dropped. Operators tried to interlock Pbar, but the interlocks wouldn't make up. An operator called in an interlock expert.

Operations established beam to NuMI at 2:39 PM. MiniBooNE beam followed a few minutes later. A problem with a Meson quadrupole power supply (MC2Q1) stopped MCenter beam. Operations called in a power supply expert.

The interlock technician found and fixed a bad relay for the Pbar line at 4:02 PM.

Operations reported that the tech had repaired the Meson power supply and that Operations had resumed sending beam to MCenter by 5:15 PM.

Pbar resumed stacking at 6:09 PM.

At 7:35 PM, Operations reported that MC2Q1 was tripping off again. At 11:19 PM, it tripped off and wouldn't reset. Since MCenter had already scheduled a weeklong shutdown to begin on Monday, they decided to begin their shutdown early.

Monday June 6
The midnight shift began with Operations monitoring store 4193 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

Operations terminated store 4193 at 1 AM.

Operations put the TeV into shot setup at 1:15 AM.

NuMI resumed taking beam at 3:03 AM.

Operations established store 4194 at 3:20 AM with an initial luminosity of 87.3E30. Pbar resumed stacking.

The Plan for Monday
The plans for today are to monitor the store and to stack and stash. During the Tuesday midnight shift, Operations will establish a new combination store; stashing will halt, but stacking will continue. The Main Injector and the Booster will be accessed for maintenance work at 8 AM Tuesday morning, as long as there is a store in the TeV.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/6/2005   email Fermilab